# Unified People Tracking with Graph Neural Networks

### v1

@@ 1 Introduction
Multi-People Tracking is a well-researched topic where tracking-by-detection has become the de facto standard. Algorithms relying on Intersection-over-Union (IoU) matching to associate detections into trajectories (Zhang et al., 2022; Bergmann et al., 2019; Wojke et al., 2017) currently top state-of-the-art benchmarks. While performance has steadily improved due to the availability of high-quality detectors, association algorithms remain largely handcrafted and highly specialized. For example, standard monocular association strategies are not equipped to handle multi-view data, and vice versa.
Single-view models often neglect to exploit the underlying scene structure and only focus on appearance and bounding-box information. Multi-view methods incorporate scene structure but often only at detection time to remove the need for association across views. Moreover, existing tracking benchmarks focus on short sequences with minimal scene occlusions. In contrast, real-world applications require handling long sequences with severe occlusions caused by people occluding each other and environmental structures, such as walls and columns, hiding them. These occlusions can cause extended disappearances and pose significant challenges to current tracking algorithms.
In this work, we introduce a unified, fully differentiable model that replaces traditional, manually designed association strategies. It learns to associate detections directly into trajectories without relying on pre-computed tracklets. Our approach builds a spatiotemporal graph where edges connect detections both temporally and across camera views in multi-view scenarios. By simultaneously aggregating appearance, motion, and contextual cues through message passing in the graph, our model enables seamless information flow across entire sequences. The graph structure is dynamically updated during both training and inference to handle streaming data in an online fashion. Our model assigns probabilities to the graph’s edges and vertices to indicate whether they belong to the same person’s trajectory. These probabilities are then used by an efficient online algorithm to extract the most likely trajectories.
Due to the limited scale of the existing datasets and their lack of scene structure information, we created our own. This new dataset is the largest of its kind, containing 25 partially overlapping views and more than 400 meters of tracked paths. It comprises two sequences of 20 minutes each, equivalent to 24,000 frames per camera at 10 FPS. Additionally, it provides camera calibrations and detailed scene mesh information. The dataset can be used to train and evaluate models in both single-view and multi-view scenarios, as well as to develop new methods to leverage scene structure for tracking.
@@
    State-of-the-art performance: The only method achieving competitive results on both single-view and multi-view people tracking benchmarks without sequence-specific tuning.
Extensive experimentation shows that our approach matches the state-of-the-art in single-camera setups and far exceeds it in multi-camera ones. This is significant because most realistic surveillance scenarios involve multiple cameras, some of which have overlapping fields of views and some not. Thus, a method that does well in both cases is what is needed in practice. Our dataset showcases this in a real-world environment. Both our dataset and code are publicly available to foster further research [^1] Dataset: https://scout.epfl.ch/, Code: https://github.com/cvlab-epfl/UMPN.
@@ 2 Related Work
Multi-target tracking is a core computer vision task with applications to surveillance, autonomous driving, medical imaging, and retail automation (Ciaparrone et al., 2020; Yilmaz et al., 2006; Smeulders et al., 2014). Most current methods rely on tracking-by-detection (Andriluka et al., 2008), where targets detected in consecutive frames are associated to produce complete trajectories (Felzenszwalb et al., 2010; Ren et al., 2015; Girshick, 2015; Zhou et al., 2019). Despite progress, challenges remain in crowded scenes with occlusions.
@@
Traditional approaches use handcrafted association methods, including probabilistic filters like JPDA (Bar-Shalom & Fortmann, 1988) and MHT (Reid, 1979) for motion prediction, and combinatorial optimization techniques such as dynamic programming (Fleuret et al., 2008). Graph-based formulations, where nodes represent detections or fixed spatial locations, have been widely studied using k-shortest paths (Berclaz et al., 2011), network flow (BenShitrit et al., 2014; Wang et al., 2019), and multi-cuts (Tang et al., 2017; 2016). While efficient, they rely on hand-designed objective functions and affinity metrics and do not generalize well to multi-view association.
@@
Learning-based methods estimate association probabilities between detections. Early approaches used conditional random fields (Yang & Nevatia, 2012) or deep networks like CNNs (Son et al., 2017) and RNNs (Sadeghian et al., 2017; Milan et al., 2017) for pairwise affinities. Recent methods (He et al., 2021; Brasó & Leal-Taixé, 2020; Kim et al., 2022; Li et al., 2020) leverage graph neural networks (Kipf & Welling, 2016; Gilmer et al., 2017) to learn affinities directly on detection graphs, while others (Wojke et al., 2017; Ristani & Tomasi, 2018; Lv et al., 2024; Qin et al., 2024) focus on learning discriminative appearance features. However, most GNN-based trackers (Brasó & Leal-Taixé, 2020; Cetintas et al., 2023) operate offline on static graphs, leveraging future detections for association decisions. In contrast, our model is designed to make fully online decisions based only on past detections.
@@
Multi-view tracking provides additional robustness through geometric constraints while introducing challenges in data fusion (Kamal et al., 2013). Some methods first perform per-view tracking before cross-view association (Bredereck et al., 2012; Lan et al., 2020; Nguyen et al., 2021), while others leverage geometry by estimating occupancy maps (Berclaz et al., 2011; Baqué et al., 2017) or projecting features onto a common ground plane (Hou et al., 2020; Song et al., 2021; Engilberge et al., 2023a). Recent approaches (Ong et al., 2020; Cheng et al., 2023b) explore associating monocular detections across both views and time, though they use a two-step process preventing joint reasoning. Our unified model simultaneously integrates inter-view, contextual and temporal cues.
@@
Early datasets like PETS (Ferryman & Shahrokni, 2009a) and TUD (Andriluka et al., 2008) focused on simple scenarios. Modern benchmarks like MOT Challenge (Milan et al., 2016; Dendorfer et al., 2021) provide diverse crowded sequences with standardized evaluation protocols. For multi-view scenarios, datasets like WILDTRACK (Chavdarova et al., 2018) and Campus (Fleuret et al., 2008) offer synchronized multi-camera sequences with calibration. While SCOUT has a lower density than MOT20, its challenges are of a different nature with significantly longer trajectories (on average 2.8x longer) and frequent multi-second occlusions behind structural elements. Thanks to multi-view annotation, we can track objects as they pass behind columns and walls—scenarios extremely challenging for single-view methods.
@@ 3 Approach
We formulate the people tracking problem as an edge classification task on a weighted spatio-temporal graph connecting object detections. Edge classification assigns probabilities of an edge being active, meaning that the two detections it links belong to the same person. This being done, we find optimal trajectories in the graph.
@@ 3.1 Spatio-Temporal Graph
Given all the detections in all images $d_{i}^{t,j}$, we build a spatio-temporal graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ depicted Fig. 1. Its vertices $\mathcal{V}=\{d_{i}^{t,j}\}$ are the detections. They are linked by a set of edges $\mathcal{E}$ of three different kinds:
@@
The graph starts empty and detections are added in a streaming fashion each time a new image is acquired, until it covers a window of $W$ time-steps, where $W$ is a model parameter. Then, the weighting algorithm described below classifies edges linking the same individuals. When a new frame with new detections becomes available, the temporal window is moved recursively by one time step, discarding the nodes and edges connected to the oldest frame in the graph, adding new detections, and creating new edges. The algorithms runs again on the updated graph. This is illustrated at the bottom of Fig. 2. In practice, we connect all detections within a fixed spatial and temporal distance. We provide more details in Section 5.
@@
With the growing interest in embodied AI, the need for scene understanding is becoming more and more important. In such contexts, knowledge about the scene is often available in the form of point clouds, meshes, or even semantic maps. We can use the graph structure to encode this knowledge by introducing camera vertices and edges. The camera vertices are connected to the detection vertices with camera edges. The feature vectors of the camera vertex are initialized using the camera parameters. The camera edges are used to encode the prior knowledge about the scene. They are initialized using the distance between the camera and the detection, and the angular distance between the camera line of sight and the detection location in the world coordinates.
@@ 3.2 Weighting the Edges using a GNN
$\mathbf{v}^{0}$ $=[\text{MLP}\left(\mathbf{p}_{d}\right)|\text{MLP}\left(\mathbf{w}_{d}\right)|\text{MLP}\left(\mathbf{t}_{d}\right)]\;$
@@
The edge representations $\mathbf{e}$ are updated first. Each $\mathbf{e}$ is concatenated with the representations $\mathbf{v}$ of the vertices it connects. This is then fed to a two-layer MLP with layer normalization (Ba et al., 2016) and ReLU activation (Fukushima, 1975) between them. In what follows, all MLP models introduced share this two-layer architecture unless otherwise stated.
The output of the MLP is added to the input edge representation, as in a Resnet (Szegedy et al., 2017). Formally, for an edge $e_{ij}$ connecting the vertices $v_{i}$ and $v_{j}$, the updated edge representation $\mathbf{e}_{ij}$ can be written as
@@
As for the edge update, we use three different MLPs parameterized by $\Theta^{m^{v}}$, $\Theta^{m^{t}}$, and $\Theta^{m^{c}}$ to generate the messages. Then, each message type—derived from its edge type—is aggregated by taking their for vertex $v_{i}$
@@ 3.3 Training
Given an input graph $\mathcal{G}$, one forward pass through the proposed network yields $P(\hat{\mathcal{E}}),P(\hat{\mathcal{V}}),\hat{\mathcal{G}}=\text{UMPN}(\mathcal{G},\Theta_{enc},\Theta_{u},\Theta_{\text{cls}})$. Here, $P(\hat{\mathcal{E}})$ represents the probabilities of all the temporal and inter-view edges, $P(\hat{\mathcal{V}})$ the vertex probabilities, and $\hat{\mathcal{G}}$ the graph with updated vertex and edge representations, as depicted in Fig. 2.
@@
We train our UMPN to learn the parameters $\Theta_{enc}$, $\Theta_{u}$ and $\Theta_{cls}$ so that the view and temporal edges, along the vertices, of $\mathcal{G}$ are correctly classified. To this end, we use ground-truth edge/vertex labels. Each edge $e_{i,j}$ and vertex $v_{i}$ is annotated with its ground-truth binary class label $e_{i,j}^{*}$ and $v_{i}^{*}$, respectively. We minimize a multi-task loss based on focal losses (Lin et al., 2017). We write it as
@@
$\mathcal{L}$ (${}_{\text{focal}}(y,p)=-y(1-p)^{\gamma}\log(p)-(1-y)p^{\gamma}\log(1-p)\;,$ (8)
where $\mathcal{L}_{\text{focal}}$ is the focal loss for a binary classification task, $y\in\{0,1\}$ is the ground truth label, $p\in[0,1]$ is the prediction, and $\gamma$ is a focusing parameter. This provides a robust loss that down-weights the contribution of easy examples.
@@
Recall from Section 3.1, that we recursively build spatio-temporal graphs that cover temporal windows of length $W$. Each time it is updated, we compute the supervised training loss of Eq. 8. We then move the temporal window by one time step, discarding the oldest nodes and edges from time step $t-W+1$ and adding new ones from time step $t+1$. In the resulting graph, the node and edge representations between $t-W+2$ and $t+1$ are the results of the previous iteration and already encode spatio-temporal context. We iterate this process until the end of the sequence. As detailed in Section 5.1, the loss is accumulated over several frames before being backpropagated to update the model parameters.
@@ 4 Datasets
Dataset #Cam. FPS #Frames #IDs Geom.
MOT17 [41] 1 14-30 11k 1.3k None
MOT20 [16] 1 25 13k 3.5k None
PETS2009 [21] 7 7 795 19 Plane
WILDTRACK [12] 7 2 2.8k 313 Plane
SCOUT (Ours) 25 10 564k 3k Mesh
@@
We evaluate our approach on three publicly available datasets, WILDTRACK (Chavdarova et al., 2018), MOT17 (Milan et al., 2016) and MOT20 (Dendorfer et al., 2021). We use the MOT17 train/val split from (Zhou et al., 2020). As these datasets are relatively small and do not provide much information about the structure of the surrounding scenes, we created our own larger-scale one that covers both single- and multi-view scenarios. As shown in Table 1, it features 25 calibrated cameras and 564,000 frames, which is much more than the other two. It includes camera calibration and 3D scene geometry for both single and multi-view evaluation. We dub it SCOUT for Scene Context and Occlusion Understanding. In the remainder of this section, we explain how it was created.
@@ 4.1 Data Acquisition
In Fig. 3, we show the placement and the region of interest for each camera.
@@
The collection was run for two windows of 1h each. From each window, sequences of 20 minutes were selected and sub-sampled at 10 Hz frame rate, which resulted in 12000 synchronized frames per sequence and per camera. Due to faults during the acquisition campaign, one of the sequences features only 22 cameras while all the others feature 25.
@@ 4.2 Scene Reconstruction
We provide the scene reconstruction as a textured mesh in a metric coordinate system. We carried a 360 degree camera equipped with a GPS receiver through the entire scene, overlapping with the fields of view of all the cameras. The mesh was then built using OpenSFM (Mapillary, 2021). Fig. 4 illustrates sample scene reconstructions.
@@
The cameras are calibrated using a pinhole model (Tsai, 1987). While the intrinsic parameters are computed using a calibration target, the extrinsics are obtained using the scene reconstruction as an intermediate step. More specifically, we run a PnP algorithm (Lepetit et al., 2009) on the dense matches between the 360 degree footage and the individual camera frames to estimate the camera poses.

### v2

# One Graph to Track Them All: Dynamic GNNs for Single- and Multi-View Tracking

@@ 1 Introduction
Multi-People Tracking is a well-researched topic where tracking-by-detection has become the de facto standard. Algorithms relying on Intersection-over-Union (IoU) matching to associate detections into trajectories (Zhang22d; Bergmann19a; Wojke17) currently top state-of-the-art benchmarks. More recently, learning-based approaches such as transformer-based methods ( (Zeng22b; Meinhardt22)) and graph neural networks ( (Cetintas23)) have been proposed. While performance has steadily improved due to the availability of high-quality detectors, association algorithms remain highly specialized. For example, standard monocular association strategies are not equipped to handle multi-view data, and vice versa.
Single-view models often neglect to exploit the underlying scene structure and only focus on appearance and bounding-box information. Multi-view methods incorporate scene structure but often only at detection time to remove the need for association across views. Moreover, existing tracking benchmarks focus on short sequences with minimal scene occlusions. In contrast, real-world applications require handling long sequences with severe occlusions caused by people occluding each other and environmental structures, such as walls and columns, hiding them. These occlusions can cause extended disappearances and pose significant challenges to current tracking algorithms. While existing benchmarks feature lots of occlusions caused by people hiding each other, scene-induced occlusions remains underexplored.
In this work, we introduce a unified, fully differentiable model that replaces traditional association strategies. It learns to associate detections directly into trajectories in an online manner. Our approach builds a spatiotemporal graph where edges connect detections both temporally and across camera views in multi-view scenarios. By simultaneously aggregating appearance, motion, and contextual cues through message passing in the graph, our model enables seamless information flow across entire sequences. The graph structure is dynamically updated during both training and inference to handle streaming data in an online fashion. Our model assigns probabilities to the graph’s edges and vertices to indicate whether they belong to the same person’s trajectory. These probabilities are then used by an efficient online algorithm to extract the most likely trajectories.
Since existing datasets lack information about scene structure, we created our own. This new dataset is the largest of its kind, containing 25 partially overlapping views and more than 400 meters of tracked paths. It comprises two sequences of 20 minutes each, equivalent to 24,000 frames per camera at 10 FPS. Additionally, it provides camera calibrations and detailed scene mesh information. The dataset can be used to train and evaluate models in both single-view and multi-view scenarios, as well as to develop new methods to leverage scene structure for tracking.
@@
    State-of-the-art performance: The only method achieving competitive results on both single-view and multi-view people tracking benchmarks.
Extensive experimentation shows that our approach matches the state-of-the-art in single-camera setups and far exceeds it in multi-camera ones. This is significant because most realistic surveillance scenarios involve multiple cameras, some of which have overlapping fields of views and some not. Thus, a method that does well in both cases is what is needed in practice. Our dataset showcases this in a real-world environment. Both our dataset and code are publicly available to foster further research.
 [^1] Dataset: https://scout.epfl.ch/, Code: https://github.com/cvlab-epfl/UMPN
@@ 2 Related Work
Multi-target tracking is a core computer vision task with applications to surveillance, autonomous driving, medical imaging, and retail automation (Ciaparrone20; Yilmaz06; Smeulders14). Most current methods rely on tracking-by-detection (Andriluka08), where targets detected in consecutive frames are associated to produce complete trajectories (Felzenszwalb10; Ren15; Girshick15; Zhou19a). Despite progress, challenges remain in crowded scenes with occlusions.
@@
Traditional approaches use handcrafted association methods, including probabilistic filters like JPDA (BarShalom88) and MHT (Reid79) for motion prediction, and combinatorial optimization techniques such as dynamic programming (Fleuret08a). Graph-based formulations, where nodes represent detections or fixed spatial locations, have been widely studied using k-shortest paths (Berclaz11), network flow (BenShitrit14; Wang19f), and multi-cuts (Tang17; Tang16). While efficient, they rely on hand-designed objective functions and affinity metrics and do not generalize well to multi-view association.
@@
Learning-based methods estimate association probabilities between detections. Early approaches used conditional random fields (Yang12a) or deep networks like CNNs (Son17) and RNNs (Sadeghian17; Milan17) for pairwise affinities. Recent methods (He21c; Braso20; Kim22; li20m) leverage graph neural networks (Kipf16; Gilmer17) to learn affinities directly on detection graphs, while others (Wojke17; Ristani18; Lv24; Qin24b) focus on learning discriminative appearance features. However, most GNN-based trackers (Braso20; Cetintas23) operate offline on static graphs, leveraging future detections for association decisions. In contrast, our model is designed to make fully online decisions based only on past detections.
@@
Multi-view tracking provides additional robustness through geometric constraints while introducing challenges in data fusion (Kamal13). Some methods first perform per-view tracking before cross-view association (Bredereck12; Lan20; Nguyen21), while others leverage geometry by estimating occupancy maps (Berclaz11; Baque17b) or projecting features onto a common ground plane (Hou20a; Song21e; Engilberge23a). Recent approaches (Ong20; Cheng23c) explore associating monocular detections across both views and time, though they use a two-step process preventing joint reasoning. Our unified model simultaneously integrates inter-view, contextual and temporal cues.
@@
Early datasets like PETS (Ferryman09a) and TUD (Andriluka08) focused on simple scenarios. Modern benchmarks like MOT Challenge (Milan16b; Dendorfer21) provide diverse crowded sequences with standardized evaluation protocols. For multi-view scenarios, datasets like WILDTRACK (Chavdarova18a) and Campus (Fleuret08a) offer synchronized multi-camera sequences with calibration. While SCOUT has a lower density than MOT20, its challenges are of a different nature with significantly longer trajectories (on average 2.8x longer) and frequent multi-second occlusions behind structural elements. Thanks to multi-view annotation, we can track objects as they pass behind columns and walls—scenarios extremely challenging for single-view methods.
@@ 3 Approach
We took our inspiration from the many classical method that rely on constructing a graph to track people. While most previous graph-based methods were designed for offline tracking in either single- or multi-view settings, our approach is fully online and generalizes to both scenarios. As in (Berclaz11; Braso20; Zhang08a; Wang19f), we formulate the tracking problem as an edge classification task in a weighted spatio-temporal graph connecting object detections. Edge classification assigns probabilities of an edge being active, meaning that the two detections it links belong to the same person. This being done, we find optimal trajectories in the graph.
@@ 3.1 Spatio-Temporal Graph
Given all the detections in all images $d_{i}^{t,j}$, we build a spatio-temporal graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ depicted Fig.˜1. Its vertices $\mathcal{V}=\{d_{i}^{t,j}\}$ are the detections. They are linked by a set of edges $\mathcal{E}$ of three different kinds:
@@
The graph starts empty and detections are added in a streaming fashion each time a new image is acquired, until it covers a window of $W$ time-steps, where $W$ is a model parameter. Then, the weighting algorithm described below classifies edges linking the same individuals. When a new frame with new detections becomes available, the temporal window is moved recursively by one time step, discarding the nodes and edges connected to the oldest frame in the graph, adding new detections, and creating new edges. The algorithms runs again on the updated graph. This is illustrated at the bottom of Fig.˜2. In practice, we connect all detections within a fixed spatial and temporal distance. We provide more details in Section˜5.
@@
ă With the growing interest in embodied AI, the need for scene understanding is becoming more and more important. In such contexts, knowledge about the scene is often available in the form of point clouds, meshes, or even semantic maps. We can use the graph structure to encode this knowledge by introducing camera vertices and edges. The camera vertices are connected to the detection vertices with camera edges. The feature vectors of the camera vertex are initialized using the camera parameters. The camera edges are used to encode the prior knowledge about the scene. They are initialized using the distance between the camera and the detection, and the angular distance between the camera line of sight and the detection location in the world coordinates.
@@ 3.2 Weighting the Edges using a GNN
$\mathbf{v}^{0}$ $=[\text{MLP}_{p}\left(\mathbf{p}_{d}\right)|\text{MLP}_{w}\left(\mathbf{w}_{d}\right)|\text{MLP}_{t}\left(\mathbf{t}_{d}\right)]\;$
@@
The edge representations $\mathbf{e}$ are updated first. Each $\mathbf{e}$ is concatenated with the representations $\mathbf{v}$ of the vertices it connects. This is then fed to a two-layer MLP with layer normalization (Ba16) and ReLU activation (Fukushima75) between them. In what follows, all MLP models introduced share this two-layer architecture unless otherwise stated.
The output of the MLP is added to the input edge representation, as in a Resnet (Szegedy17). Formally, for an edge $e_{ij}$ connecting the vertices $v_{i}$ and $v_{j}$, the updated edge representation $\mathbf{e}_{ij}$ can be written as
@@
As for the edge update, we use three different MLPs parameterized by $\Theta^{m^{v}}$, $\Theta^{m^{t}}$, and $\Theta^{m^{c}}$ to generate the messages. Then, each message type—derived from its edge type—is aggregated by taking their vertex $v_{i}$
@@ 3.3 Training
[…]
@@
We train our UMPN to learn the parameters $\Theta_{enc}$, $\Theta_{u}$ and $\Theta_{cls}$ so that the view and temporal edges, along the vertices, of $\mathcal{G}$ are correctly classified. To this end, we use ground-truth edge/vertex labels. Each edge $e_{i,j}$ and vertex $v_{i}$ is annotated with its ground-truth binary class label $e_{i,j}^{*}$ and $v_{i}^{*}$, respectively. We minimize a multi-task loss based on focal losses (Lin17f). We write it as
@@
$\mathcal{L}$ $_{\text{focal}}(y,p)=-y(1-p)^{\gamma}\log(p)-(1-y)p^{\gamma}\log(1-p)\;,$ (8)
ă where $\mathcal{L}_{\text{focal}}$ is the focal loss for a binary classification task, $y\in\{0,1\}$ is the ground truth label, $p\in[0,1]$ is the prediction, and $\gamma$ is a focusing parameter. This provides a robust loss that down-weights the contribution of easy examples.
@@
Recall from Section 3.1, that we recursively build spatio-temporal graphs that cover temporal windows of length $W$. Each time it is updated, we compute the supervised training loss of Eq.˜8. We then move the temporal window by one time step, discarding the oldest nodes and edges from time step tW+1t-W+1 and adding new ones from time step $t+1$. In the resulting graph, the node and edge representations between tW+2t-W+2 and $t+1$ are the results of the previous iteration and already encode spatio-temporal context. We iterate this process until the end of the sequence. As detailed in Section 5.1, the loss is accumulated over several frames before being backpropagated to update the model parameters.
@@ 4 Datasets
Dataset #Cam. FPS #Frames #IDs Geom. Avg. Clip Length #Clips
MOT17 [Milan16b] 1 14-30 11k 1.3k None 33s 14
MOT20 [Dendorfer21] 1 25 13k 3.5k None 67s 8
PETS2009 [Ferryman09a] 7 7 795 19 Plane 22s 4
WILDTRACK [Chavdarova18a] 7 2 2.8k 313 Plane 3min20s 1
SCOUT (Ours) 25 10 564k 3k Mesh 20min 2
@@
We evaluate our approach on three publicly available datasets, WILDTRACK (Chavdarova18a), MOT17 (Milan16b) and MOT20 (Dendorfer21). We use the MOT17 train/val split from (Zhou20b). As these datasets are relatively small and do not provide much information about the structure of the surrounding scenes, we created our own larger-scale one that covers both single- and multi-view scenarios. As shown in Table˜1, it features 25 calibrated cameras and 564,000 frames, which is much more than the other two. It includes camera calibration and 3D scene geometry for both single and multi-view evaluation. We dub it SCOUT for Scene Context and Occlusion Understanding. In the remainder of this section, we explain how it was created.
@@ 4.1 Data Acquisition
In Fig.˜3, we show the placement and the region of interest for each camera.
@@
The collection was run for two windows of 1h each. From each window, sequences of 20 minutes were selected and sub-sampled at 10 Hz frame rate, which resulted in 12000 synchronized frames per sequence and per camera. Due to faults during the acquisition campaign, one of the sequences features only 22 cameras while the other one features 25.
@@ 4.2 Scene Reconstruction
We provide the scene reconstruction as a textured mesh in a metric coordinate system. We carried a 360 degree camera equipped with a GPS receiver through the entire scene, overlapping with the fields of view of all the cameras. The mesh was then built using OpenSFM (OpenSFM). Fig.˜4 illustrates sample scene reconstructions.
@@
The cameras are calibrated using a pinhole model (Tsai87). While the intrinsic parameters are computed using a calibration target, the extrinsics are obtained using the scene reconstruction as an intermediate step. More specifically, we run a PnP algorithm (Lepetit09) on the dense matches between the 360 degree footage and the individual camera frames to estimate the camera poses.